OpenStudy (anonymous):

1) Evaluate lim x->1 (x-1/Sqrt(x)-1) 2)Evaluate limx->0 (1-e^(2x)/1-e^x)

OpenStudy (callisto):

\[lim_{x \to 1} \frac{x-1}{\sqrt{x}-1}\]For the first one?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

find the first derivative of numerator and denominator and put value of limit x -> 1 your will get your answer

OpenStudy (callisto):

Hmm.. multiply a conjugate\[ \lim_{x \to 1} \frac{x-1}{\sqrt{x}-1}\]\[=\lim_{x \to 1} \frac{x-1}{\sqrt{x}-1}\times \frac{\sqrt{x}+1}{\sqrt{x}+1}\]\[=\lim_{x \to 1} \frac{(x-1)(\sqrt{x}+1)}{(\sqrt{x}-1)(\sqrt{x}+1)}\]Expand the bottom part.

OpenStudy (hba):

I guess it tends to 2. (x-1)=(sqrt{x}-1)(sqrt{x}+1}

OpenStudy (callisto):

Use \(a^2 - b^2 = (a+b)(a-b)\) to expand it. Cancel what you can cancel

OpenStudy (hba):

You do not have to do the conjugate see this....\[\lim_{x \to 1} \frac{x-1}{\sqrt{x}-1}\] \[\lim_{x \rightarrow 1}\frac{ (\sqrt{x}-1)(\sqrt{x}+1) }{ \sqrt{x}-1 }\]

OpenStudy (goten77):

cant you just use l hospitals rule on problem 1?

OpenStudy (hba):

No

OpenStudy (hba):

Now cancel the terms from the numerator and denominator @Tech-Hunter

OpenStudy (hba):

Then apply limit and kaboom you get your answer.

OpenStudy (goten77):